THE ‘4050 Å GROUP’ OF THE Ã^{1}Π_{u} - X̃^{1}Σ_{g}^{+} TRANSITION SYSTEM OF C_{3}

\underline{D. ZHAO}, H. LINNARTZ, Sackler Laboratory for Astrophysics, Leiden Observatory, University of Leiden, P.O. Box 9513, NL-2300 RA Leiden, the Netherlands; M. A. HADDAD, W. UBACHS, Laser-LaB, VU University Amsterdam, De Boelelaan 1081, NL-1081 HV, Amsterdam, The Netherlands; M. R. SCHMIDT, Department of Astrophysics, N. Copernicus Astronomical Center, ul. Rabiańska 8, 87-100 Toruń, Poland; J. KRELOWSKI, Centre for Astronomy, Nicolaus Copernicus University, Gagarina 11, 87-100 Toruń, Poland; G. A. GALAZUTDINOV, Instituto de Astronomia, Universidad Catolica del Norte, Av. Angamos 0610, Antofagasta, Chile.

The ‘4050 Å group’ of C_{3}, which consists of a series of bands of the Ã^{1}Π_{u} - X̃^{1}Σ_{g}^{+} electronic transition system, has been of much interest to astronomers as it serves as a remote diagnostic of the physical-chemical conditions in translucent interstellar clouds. In this contribution, we present:

a) The high resolution (R = 80 000) detection of eight vibronic bands of C_{3} in the diffuse translucent cloud towards HD 169454. Four of these vibronic bands are also detected in two additional objects: HD 73882 and HD 154368. Column densities and excitation conditions of C_{3} are inferred.

b) A laboratory re-examination of the eight vibronic bands observed towards HD 169454 using cavity ring-down spectroscopy and a supersonic plasma jet. High-quality laboratory data build the list of lines with a wavelength accuracy of <0.01Å. This is sufficient for the analysis of the observational data. Improved spectroscopic parameters of the corresponding vibronic states are presented. An improved perturbation analysis in the Ã^{1}Π_{u} (000) state is given as well.

c) The systematic laboratory investigation on the Ã^{1}Π_{u} - X̃^{1}Σ_{g}^{+} 000-000 electronic origin band of ^{13}C-substituted C_{3}. Rotationally resolved spectra of all five ^{13}C-isotopologues are recorded in a supersonic plasma expansion by discharging 13C_{2}H_{2} or ^{12}C_{2}H_{2}/^{13}C_{2}H_{2} mixtures diluted in noble gas. The Ã^{1}Π_{u} state molecular constants for all five isotopologues and ground-state molecular constants for ^{13}C^{12}C^{13}C and ^{12}C^{13}C^{13}C are experimentally determined for the first time. This work extends the recent mid-infrared work on the ^{13}C-isotopologues of C_{3} by Giesen et al.^{a}

aT. Giesen, private communication, December 2012.